OpenStudy (anonymous):

A farmer is building a pentagon(that is 5-sided) pen to hold his animals. each side has an equal length. What is the total perimiter of the chicken pen if the length of one side is 3√5 yards?(recall that the perimiter of each side of a shape or boundary.

OpenStudy (anonymous):

Perimeter is the sum of all the sides, or 5 times the length of one side.

OpenStudy (anonymous):

possible answers are 8√5 or 15√5 or 8√25 or 15√25

OpenStudy (anonymous):

\[5 \times 3\sqrt{5} = 15\sqrt{5}\]

OpenStudy (anonymous):

Make sure you understand how to multiply this stuff though... don't just grab the answer and go

OpenStudy (anonymous):

Does that multiply expression make sense?

OpenStudy (anonymous):

If the sides of the pen were 4 feet long, the perimeter would be 5 x 4 = 20 feet

OpenStudy (anonymous):

some what could u make an example problem

OpenStudy (anonymous):

You can add and multiply things with square roots in them sort of like they were variables like x or y or whatever. You can add any two terms that have the SAME thing in the square root. Like this: \[2\sqrt{5} + 3\sqrt{5} = 5\]

OpenStudy (anonymous):

just like 2x + 3x = 5x

OpenStudy (anonymous):

In the same way, you can multiply a number times some square root term, just like it was a variable... so if 3 times 2x equals 6x, you could also have :\[3\times2\sqrt{5}=6\sqrt{5}\]

OpenStudy (anonymous):

oh ok now what about this problem A father wants to build an octoganal (that is 8-sided) play pen for his toddler . Each side of the octogon has a length of 3√50 inches. What is the perimiter of the octogon ?

OpenStudy (anonymous):

You tell me... :) Take a shot at it...

OpenStudy (anonymous):

pretend "square root 50" is called "x" so a side is 3x inches long

OpenStudy (anonymous):

So um hold up let me think

OpenStudy (anonymous):

how of ten do u get on here

OpenStudy (anonymous):

Some :) Just depends.. I try to check daily, but I can't be on all day every day

OpenStudy (anonymous):

If that "x" thing is confusing, just forget it... some people panic seeing the square roots, but if the "x" makes it worse, just go back to my examples earlier.

OpenStudy (anonymous):

im in a recovery class and im trying to make up my credits so if u help me i can send you 50 dollors at the end of every week. You are my only hope in me being able to walk across the stage and recive my diploma so what do u say can u help me out

OpenStudy (anonymous):

please

OpenStudy (anonymous):

I can't do that... But I'm not your only hope... there are a bunch of good folks on here daily. You don't have to find me personally... just ask the questions and someone will help.

OpenStudy (anonymous):

I'm happy to help, no charge, while I'm online though :)

OpenStudy (anonymous):

so would the answer to the last problem be 3√400

OpenStudy (anonymous):

Good try, but you multiply the 8 (# of sides) times the 3 in front of the square root, not times the 50 inside the square root. \[8 \times3\sqrt{50}= 24\sqrt{50}\]

OpenStudy (anonymous):

24√50 is not on the answer sheet

OpenStudy (anonymous):

24√400

OpenStudy (anonymous):

Hm... what else is on there?

OpenStudy (anonymous):

Maybe there would be a answer \[96\sqrt{2}\]

OpenStudy (anonymous):

Wait, you can simplify... sorry I missed this earlier...\[8 \times3\sqrt{50} = 24\sqrt{50} = 24\sqrt{(25)(2)}=(24)(5)\sqrt{2}=120\sqrt{2}\]

OpenStudy (anonymous):

@MBSHAW95 update to our earlier perimeter problem... I missed a simplification